OpenStudy (anonymous):

1. I am completely missing something here plz help: What is the result of subtracting 1 1/2 from the sum of 4 2/3 and 5 2/5? 2. Evaluate. 3 1/6 + 8 2/9 - 1 1/2

OpenStudy (calculusxy):

Easier way is to put in an equation: \[(4\frac{ 2 }{ 3 }-5\frac{ 2 }{ 5 })-1\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

I thought "and" meant to multiply.

OpenStudy (calculusxy):

I couldn't quite get you there.

OpenStudy (anonymous):

Well theres two questions. 1. What is the result of subtracting 1 1/2 from the sum of 4 2/3 and 5 2/5? 2. Evaluate. 3 1/6 + 8 2/9 - 1 1/2

OpenStudy (tkhunny):

"the sum of A and B" means A+B "the product of A and B" means A*B A sum needs two arguments. It's just listing them.

OpenStudy (calculusxy):

I don't really get your problem. Are you supposed to add the result of A and B or find the product of A and B?

OpenStudy (anonymous):

The questions on the page is exactly how I typed them. I'm trying to help my son with his 5th grade math homework, and for some reason, the answer I am coming up with is not right or it is and I'm not certain. Its been a very long time since I done this kind of math.

OpenStudy (calculusxy):

Okay i will try to do questions 1 and 2. lets see if we match or not.

OpenStudy (anonymous):

ty so very much

OpenStudy (calculusxy):

 Sorry for my bad handwriting. I just couldn't use the mouse to properly write.